# Improved Treatment of Bosonic Dark Matter Dynamics in Neutron Stars: Consequences and Constraints

### v1

@@ 3.1 Case I : Non-annihilating dark matter
To illustrate the effects of accumulation of DM particles, we first choose $\sigma_{\chi n}/\sigma_{g}=10^{-5}$, $m_{\chi}=10$ GeV and $\expectationvalue{\sigma v}=0$, for which DM particles settle into a BEC state. In the left panel of Fig.1, we have shown the evolution of NS temperature ($T_{NS}$) (blue solid line), the critical temperature ($T_{c}$) (red solid line) and the characteristic temperature for self-gravitation ($T_{sg}$) (black dashed line). In the absence of DM annihilation, $T_{NS}$ follows the standard cooling procedure, which initiates rapid cooling after $t=10^{6}$ years. On the other hand, $T_{c}$ increases gradually due to continuous acquisition of DM particles. Thus, $T_{NS}$ becomes smaller than $T_{c}$ around $t=8\times 10^{6}$ years, which marks the BEC formation. We note that the effect of self-gravitation remains small until the BEC formation, indicated by $T_{sg}$ being smaller than $T_{c}$ at the relevant epoch. Consequently, the BH formation is to be determined by the number of DM particles, exceeding the Chandrasekhar limit ($N_{ch}$). Thus the current choice of parameters, indicated by ($\bm{\star}$) mark in Fig.3 is excluded from the observation of an old NS.
@@
In Fig.2 we show that DM particles become self-gravitating rather than forming a BEC for parameters, $m_{\chi}=10^{5}~{}{\rm GeV}$ and $\sigma_{\chi n}/\sigma_{g}=1$, which have also been shown in Fig.LABEL:fig:scn by ($\bm{\times}$). In the left panel, $T_{NS}$ falls below $T_{sg}$ around $t=6\times 10^{6}$ years, at which $T_{c}$ remains smaller than both $T_{NS}$ and $T_{sg}$. This signifies the imminent collapse of DM particles to a BH without being in a BEC state. It is also evident from the right panel plots that $N^{0}_{\chi}$ remains vanishing when $N_{\chi}$ takes over $N_{sg}$. In this scenario, $N_{sg}$ is the deciding factor of BH formation, unlike the former case where $N_{ch}$ is the relevant quantity. To note, the BH formation via self-gravitation is favorable for heavy DM scenarios, as evident from Fig.3.
@@
The green shaded region shows the self-gravitating phase leading to the formation of BH. In this case, the constraint on $\sigma_{\chi n}$ becomes weaker due to larger number of DM particles requiring to destabilize a NS. To contrast with the former case, the upper limit on $\sigma_{\chi n}$ increases as the DM mass increases in the regime, $m_{\chi}\gtrsim 10$ TeV. In this region, BH formation is governed by $N_{sg}$, which depends on both $T_{NS}$ and $m_{\chi}$ (see Eq.16), unlike the formation of BH in the BEC case, which depends only on $m_{\chi}$. Consequently, the initial mass of the nascent BH is given by, $M_{BH,i}\sim m_{\chi}N_{sg}\propto(T_{NS}/m_{\chi})^{3/2}$, which again necessitate our dynamical analysis. Similar to the BEC case, the BH evaporation effects open up the DM parameter space but for $m_{\chi}\gtrsim 3\times 10^{7}$ GeV.
@@
In addition to NS constraints, we present the current direct detection bound from XENON1TXENON:2019gfn (for GeV to TeV DM) and CRESST-III CRESST:2019jnq (mainly for sub-GeV DM) on the $\sigma_{\chi n}-m_{\chi}$ plane, which is somewhat complementary to the current exclusion limits coming from DM capture scenarios. It is worth noting that $m_{\chi}\in[12{~{}\rm GeV},3{~{}\rm TeV}]$, DD experiments remain instrumental in constraining the DM-neutron cross-section. This is an important implication of non-annihilating bosonic DM forming a BEC state, thereby evading constraints from the NS observations often advocated to probe smaller DM-neutron cross-sections in the literature. However, for sub-GeV DM and DM with masses larger than tens of a TeV, NS observations can probe various DM scenarios with greater sensitivity compared to the DD experiments, albeit various astrophysical uncertainties.
@@
Now, let us compare the present result in the context of already existing bounds in most of the literature McDermott:2011jp ; Bramante:2013hn ; Garani:2018kkd ; Bhattacharya:2023stq ; Lu:2024kiz . In these papers, the constraints on the $\sigma_{\chi n}-m_{\chi}$ plane are derived considering BEC and non-BEC state separately with a fixed NS temperature. In our work, we dynamically determine the demarcation between a BEC and a non-BEC state given a set of DM parameters. This is one of the main difference from the earlier works. Additionally, in our analysis we check thermalization of DM particles throughout the evolution, whereas in previous studies the thermalization timescale is derived considering a fixed final temperature of a NS. However, the upper bound on $\sigma_{\chi n}$ coming from the BH formation in the BEC case, does not disagree to a great extent comapred to these works. The results for the BEC and non-BEC states from Fig.1 of Ref.McDermott:2011jp have been shown by the red dot-dashed and green dot-dashed line respectively in our Fig.3. It is apparent that the significant difference emerges for the non-BEC case. We have checked that the recent analysis in Ref.Garani:2018kkd (depicted in Fig.10) agrees with results shown in Ref.McDermott:2011jp , thereby our results differ from both of the references.
@@
In particular, there are three differences for the BH formation from a non-BEC state. Firstly, the upper bound on $\sigma_{\chi n}$ in the mass regime, $1~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$ is not applicable due to BEC formation, as implied by our analysis. Moreover, this region is not even ruled out (as claimed in Ref.McDermott:2011jp ; Garani:2018kkd ) due to BH formation. Secondly, the constraint on $\sigma_{\chi n}$ as derived from our dynamical treatment is appreciably different for $m_{\chi}>16~{}{\rm TeV}$. This stems from the fact that the DM number needed for the onset of self-gravitation is significantly higher in our analysis than in the previous studies. In fact, given the DM parameters, the temperature, $T_{NS}$ at which the self-gravitation sets in, is completely determined. For example we see for $m_{\chi}=10^{5}$ GeV, $N_{sg}=N_{\chi}$ at $T_{NS}\approx 0.005~{}{\rm MeV}$, apparent from Fig.2. In previous studies, the onset of self-gravitation is assumed at $T_{NS}=10^{-5}~{}{\rm MeV}$, which leads to the under-estimation of DM number required for the collapse. Therefore, in our analysis, the self-gravitating regime in the non-BEC case sets in earlier, which requires larger $\sigma_{\chi n}$. Finally, the nascent BH mass becomes greater in our analysis for the same DM mass, which has been reflected in the higher cut-off in the DM mass related to BH evaporation. In particular, the DM mass cut-off, beyond which there is no constraint from the stability of a NS is found to be $10$ times larger than the non-dynamic analyses.
@@ 3.2 Case II : Annihilating dark matter
It is evident from Fig.5 that DM particles neither forms a BEC nor becomes self-gravitating for the same parameters as in Fig.2. With $\sigma_{\chi n}/\sigma_{g}=1$, $T_{NS}$ is always larger than both $T_{c}$ and $T_{sg}$ , contrary to the case with $\expectationvalue{\sigma v}=0$. In particular, for annihilating DM scenario, the higher DM capture rate leads to higher temperature for old enough NS, therefore both BEC formation and BH formation are hindered eventually.
@@
The parameter region for BEC formation is greatly modified due to annihilation. This is clear by comparing the blue solid line surrounded regions in Fig.3 and the left panel of Fig.6. BEC formation depends on both $\sigma_{\chi n}$ and $\expectationvalue{\sigma v}$. Increasing $\sigma_{\chi n}$ allows more accumulation of DM particles, thus expediting the BEC formation. On the other hand, due to annihilating nature of DM, more accumulation leads to heating up the NS. The interplay between of these two effects creates the wedge-like structure within $500{~{}\rm GeV}\lesssim m_{\chi}\lesssim 10{~{}\rm TeV}$. In particular, when $\sigma_{\chi n}$ reaches above its geometric limit no further accumulation happens and it creates the vertical blue line in that region.
@@
Now, for the case of annihilating DM the BH forms within the BEC region for more restricted choices of $\sigma_{\chi n}$. In presence of annihilation smaller values of $\sigma_{\chi n}$ correspond to less number of accumulated DM particles, hindering subsequent BH formation. Note that, the maximum mass for which BH forms in the BEC state remains the same for both annihilating and non-annihilating DM. This is a consequence of BH evaporation only, which fixes the mass cut-off. Now, the BH formation from a non-BEC state takes place for $m_{\chi}>5\times 10^{5}$ GeV for $\expectationvalue{\sigma v}=10^{-41}~{}{\rm cm^{3}s^{-1}}$, significantly different from the former case depicted in Fig.3. This is because the DM abundance is reduced via annihilation, thereby requiring heavier DM particles for the onset of self-gravitation. We emphasize that for parameter points shown in Fig.3 by ($\bm{\star~{},~{}\times}$) symbols, the BH does not form in the present case, contrary to the former one.
@@
Now, we compare two panels of Fig.6 to understand the effects of different values of DM annihilation cross-section. First of all, the parameter space of BEC formation gets shrunk considerably for $\expectationvalue{\sigma v}=10^{-26}~{}{\rm cm^{3}s^{-1}}$. In particular, the minimum value of $\sigma_{\chi n}$ for BEC formation is higher than in the former scenario. This is because for the same value of $\sigma_{\chi n}$, more DM particles annihilate owing to larger $\expectationvalue{\sigma v}$. For instance, the parameter point shown by ($\bm{\star}$) in Fig.6, BEC state does not form in the right panel. Besides for DM masses ranging from $10$ MeV to $10$ TeV, there is a upper limit on $\sigma_{\chi n}$, above which BEC never forms, unlike the scenario shown in the left panel. This is resulted from the competing effects between the DM capture and annihilation, as explained earlier. The final distinction from the earlier scenario is that BH never forms for such a large annihilation rate. Consequently, the constraint on $\sigma_{\chi n}-m_{\chi}$ plane is completely relaxed from the observation of old NSs.
@@
However, DM annihilations allows us to probe $\sigma_{\chi n}$ for wide range of DM masses with the advent of modern telescopes. The thermal emissions from old NSs in the infra-red and visible frequencies can be detectable in telescopes, such as the JWST. In particular, JWST aims to detect NSs with the red-shifted surface temperatures ($T^{\infty}_{S}$) within 2000 K to 40000 K Raj:2024kjq ; Chatterjee:2022dhp , which can determine the cooling procedure of old NSs. Therefore, one can test various late-time heating scenarios related to astrophysics of NS, including the DM annihilation effects. Here, we use the sensitivity of the JWST to probe the DM-neutron cross-section, assuming the late-time heating coming only from DM annihilations. In the both panels of Fig.6, the region above the orange dot-dashed line can be probed with the detection of thermal emissions of 10 Gyr old NS having $T^{\infty}_{S}\geq 2000$ K. To remind, the surface temperature of the NS is related to $T^{f}_{NS}$, which is dependent on $m_{\chi}C_{c}$ (see Eq.14) in the capture-annihilation equilibrium. Consequently, $\sigma_{\chi n}$ is determined with mild dependence on the DM mass, $m_{\chi}$ through $\sigma_{g}$ apparent from Eq.2.
@@ 4 Outlook
Our present study is about bosonic DM particles forming either a BEC or a non-BEC state inside a NS, depending on DM observables, namely $m_{\chi}$, $\sigma_{\chi n}$ and $\expectationvalue{\sigma v}$. The constraints on DM parameters vary from the consideration of BH formation out of two distinct thermal states. However, the distinction between a BEC state and a non-BEC state, in which BH does not form, seems to be somewhat indeterminate just from the existence of an old NS and its probable late-time heating. The distinction can be possible via probing two different NS observables, namely the mass-radius relation and the tidal deformability. In particular, the presence of a DM core at the center of a NS or a DM halo in the vicinity of it affect these two quantities, testable in gravitational wave detectors as well as radio observations of pulsars. The effect of a self-interacting bosonic DM core in the NS have already been studied in Refs.Ellis:2018bkr ; Karkevandi:2021ygv ; Buras-Stubbs:2024don and references therein. In the present context, it is apparent that different thermal states of DM particles would correspond to different equation of states (EOS), leading to impact the gravitational wave (GW) spectrum LIGOScientific:2017vwq ; Annala:2017llu . Now, for a given EOS of baryons, it is possible to distinguish between a BEC state and a non-BEC state, in turn constraining the DM parameters ghosh .

### v2

@@ 3.1 Case I : Non-annihilating dark matter
To illustrate the effects of accumulation of DM particles, we first choose $\sigma_{\chi n}/\sigma_{g}=10^{-5}$, $m_{\chi}=10$ GeV and $\expectationvalue{\sigma v}=0$, for which DM particles settle into a BEC state. In the left panel of Fig.1, we have shown the evolution of NS temperature ($T_{NS}$) (blue solid line), the critical temperature ($T_{c}$) (red solid line) and the characteristic temperature for self-gravitation ($T_{sg}$) (black dashed line). In the absence of DM annihilation, $T_{NS}$ follows the standard cooling procedure, which initiates rapid cooling after $t=10^{6}$ years. On the other hand, $T_{c}$ increases gradually due to continuous acquisition of DM particles. Thus, $T_{NS}$ becomes smaller than $T_{c}$ around $t=8\times 10^{6}$ years, which marks the BEC formation. We note that the effect of self-gravitation remains small until the BEC formation, indicated by $T_{sg}$ being smaller than $T_{c}$ at the relevant epoch. Consequently, the BH formation is to be determined by the number of DM particles, exceeding the Chandrasekhar limit ($N_{ch}$). Thus the current choice of parameters, indicated by ($\bm{\star}$) mark in Fig.3 is excluded from the observation of an old NS. As an aside, we note the evolution of $T_{c}$ is mostly unaffected with a different choice of the NS mass, i.e. $M_{NS}=2.2M_{\odot}$, keeping the radius and baryon density fixed as mentioned earlier. This has been shown by the magenta dashed line in the left panel of Fig.1.
@@
In Fig.2 we show that DM particles become self-gravitating rather than forming a BEC for parameters, $m_{\chi}=10^{5}~{}{\rm GeV}$ and $\sigma_{\chi n}/\sigma_{g}=1$, which have also been shown in Fig.3 by ($\bm{\times}$). In the left panel, $T_{NS}$ falls below $T_{sg}$ around $t=6\times 10^{6}$ years, at which $T_{c}$ remains smaller than both $T_{NS}$ and $T_{sg}$. This signifies the imminent collapse of DM particles to a BH without being in a BEC state. It is also evident from the right panel plots that $N^{0}_{\chi}$ remains vanishing when $N_{\chi}$ takes over $N_{sg}$. In this scenario, $N_{sg}$ is the deciding factor of BH formation, unlike the former case where $N_{ch}$ is the relevant quantity. To note, the BH formation via self-gravitation is favorable for heavy DM scenarios, as evident from Fig.3.
@@
The green shaded region shows the self-gravitating phase leading to the formation of BH. In this case, the constraint on $\sigma_{\chi n}$ becomes weaker due to larger number of DM particles requiring to destabilize a NS. To contrast with the former case, the upper limit on $\sigma_{\chi n}$ increases as the DM mass increases in the regime, $m_{\chi}\gtrsim 10$ TeV. In this region, BH formation is governed by $N_{sg}$, which depends on both $T_{NS}$ and $m_{\chi}$ (see Eq.16), unlike the formation of BH in the BEC case, which depends only on $m_{\chi}$. Consequently, the initial mass of the nascent BH is given by, $M_{BH,i}\sim m_{\chi}N_{sg}\propto(T_{NS}/m_{\chi})^{3/2}$, which again necessitates our dynamical analysis. Similar to the BEC case, the BH evaporation effects open up the DM parameter space but for $m_{\chi}\gtrsim 3\times 10^{7}$ GeV.
@@
In addition to NS constraints, we present the current direct detection bound from XENON1TXENON:2019gfn (for GeV to TeV DM) and CRESST-III CRESST:2019jnq (mainly for sub-GeV DM) on the $\sigma_{\chi n}-m_{\chi}$ plane, which is somewhat complementary to the current exclusion limits coming from DM capture scenarios. It is worth noting that $m_{\chi}\in[12{~{}\rm GeV},3{~{}\rm TeV}]$, DD experiments remain instrumental in constraining the DM-neutron cross-section. This is an important implication of non-annihilating bosonic DM forming a BEC state, thereby evading constraints from the NS observations often advocated to probe smaller DM-neutron cross-sections in the literature. This result is quite robust with varying the local DM density, as the concerned mass range is dependent on the NS properties. In particular, the lower limit, i.e. $m_{\chi}=12$ GeV, beyond which all associated nascent BHs evaporate is fixed given the baryon density and the sound velocity inside the NS. The upper limit, i.e. $m_{\chi}=3$ TeV, beyond which BH forms due to self-gravity, is again dependent on the baryon density of the NS. However, for sub-GeV DM and DM with masses larger than tens of a TeV, NS observations can probe various DM scenarios with greater sensitivity compared to the DD experiments, albeit various uncertainties regarding NS parameters.
@@
Now, let us compare the present result in the context of already existing bounds in most of the literature McDermott:2011jp ; Bramante:2013hn ; Garani:2018kkd ; Bhattacharya:2023stq ; Lu:2024kiz . In these papers, the constraints on the $\sigma_{\chi n}-m_{\chi}$ plane are derived considering BEC and non-BEC state separately with a fixed NS temperature. In our work, we dynamically determine the demarcation between a BEC and a non-BEC state given a set of DM parameters. This is one of the main differences from the earlier works. Additionally, in our analysis we check thermalization of DM particles throughout the evolution, whereas in previous studies the thermalization timescale is derived considering a fixed final temperature of a NS. However, the upper bound on $\sigma_{\chi n}$ coming from the BH formation in the BEC case, does not disagree to a great extent comapred to these works. The results for the BEC and non-BEC states from Fig.1 of Ref.McDermott:2011jp have been shown by the red dot-dashed and green dot-dashed line respectively in our Fig.3. It is apparent that the significant difference emerges for the non-BEC case. We have checked that the recent analysis in Ref.Garani:2018kkd (depicted in Fig.10) agrees with results shown in Ref.McDermott:2011jp , thereby our results differ from both of the references.
@@
In particular, there are three differences for the BH formation from a non-BEC state. Firstly, the upper bound on $\sigma_{\chi n}$ in the mass regime, $1~{}{\rm GeV}\lesssim m_{\chi}\lesssim 5~{}{\rm TeV}$ is not applicable due to BEC formation, as implied by our analysis. Moreover, this region is not even ruled out (claimed otherwise in Ref.McDermott:2011jp ; Garani:2018kkd ) due to BH evaporation. Secondly, the constraint on $\sigma_{\chi n}$ as derived from our dynamical treatment is appreciably different for $m_{\chi}>16~{}{\rm TeV}$. This stems from the fact that the DM number needed for the onset of self-gravitation is significantly higher in our analysis than in the previous studies. In fact, given the DM parameters, the temperature, $T_{NS}$ at which the self-gravitation sets in, is completely determined. For example we see for $m_{\chi}=10^{5}$ GeV, $N_{sg}=N_{\chi}$ at $T_{NS}\approx 0.005~{}{\rm MeV}$, apparent from Fig.2. In previous studies, the onset of self-gravitation is assumed at $T_{NS}=10^{-5}~{}{\rm MeV}$, which leads to the under-estimation of DM number required for the collapse. Therefore, in our analysis, the self-gravitating regime in the non-BEC case sets in earlier, which requires larger $\sigma_{\chi n}$. Finally, the nascent BH mass becomes greater in our analysis for the same DM mass, which has been reflected in the higher cut-off in the DM mass related to BH evaporation. In particular, the DM mass cut-off, beyond which there is no constraint from the stability of a NS is found to be $10$ times larger than the non-dynamic analyses.
@@ 3.2 Case II : Annihilating dark matter
It is evident from Fig.5 that DM particles neither form a BEC nor become self-gravitating for the same parameters as in Fig.2. With $\sigma_{\chi n}/\sigma_{g}=1$, $T_{NS}$ is always larger than both $T_{c}$ and $T_{sg}$ , contrary to the case with $\expectationvalue{\sigma v}=0$. In particular, for annihilating DM scenario, the higher DM capture rate leads to higher temperature for old enough NS, therefore both BEC formation and BH formation are hindered eventually.
@@
The parameter region for BEC formation is greatly modified due to annihilation. This is clear by comparing the blue solid line surrounded regions in Fig.3 and the left panel of Fig.6. BEC formation depends on both $\sigma_{\chi n}$ and $\expectationvalue{\sigma v}$. Increasing $\sigma_{\chi n}$ allows more accumulation of DM particles, thus expediting the BEC formation. On the other hand, due to annihilating nature of DM, more accumulation leads to heating up the NS. The interplay between these two effects creates the wedge-like structure within $500{~{}\rm GeV}\lesssim m_{\chi}\lesssim 10{~{}\rm TeV}$. In particular, when $\sigma_{\chi n}$ reaches above its geometric limit no further accumulation happens and it creates the vertical blue line in that region.
@@
Now, for the case of annihilating DM the BH forms within the BEC region for more restricted choices of $\sigma_{\chi n}$. In presence of annihilation smaller values of $\sigma_{\chi n}$ correspond to less number of accumulated DM particles, hindering subsequent BH formation. Note that, the maximum mass for which BH forms in the BEC state remains the same for both annihilating and non-annihilating DM. This is a consequence of BH evaporation only, which fixes the mass cut-off. Now, the BH formation from a non-BEC state takes place for $m_{\chi}>5\times 10^{5}$ GeV for $\expectationvalue{\sigma v}=10^{-41}~{}{\rm cm^{3}~{}s^{-1}}$, significantly different from the former case depicted in Fig.3. This is because the DM abundance is reduced via annihilation, thereby requiring heavier DM particles for the onset of self-gravitation. We emphasize that for parameter points shown in Fig.3 by ($\bm{\star~{},~{}\times}$) symbols, the BH does not form in the present case, contrary to the former one.
@@
Now, we compare two panels of Fig.6 to understand the effects of different values of DM annihilation cross-section. First of all, the parameter space of BEC formation gets shrunk considerably for $\expectationvalue{\sigma v}=10^{-26}{~{}\rm cm^{3}~{}s^{-1}}$. In particular, the minimum value of $\sigma_{\chi n}$ for BEC formation is higher than in the former scenario. This is because for the same value of $\sigma_{\chi n}$, more DM particles annihilate owing to larger $\expectationvalue{\sigma v}$. For instance, the parameter point shown by ($\bm{\star}$) in Fig.6, BEC state does not form in the right panel. Besides for DM masses ranging from $10$ MeV to $10$ TeV, there is a upper limit on $\sigma_{\chi n}$, above which BEC never forms, unlike the scenario shown in the left panel. This is resulted from the competing effects between the DM capture and annihilation, as explained earlier. The final distinction from the earlier scenario is that BH never forms for such a large annihilation rate. Consequently, the constraint on $\sigma_{\chi n}-m_{\chi}$ plane is completely relaxed from the observation of old NSs.
@@
However, DM annihilations allows us to probe $\sigma_{\chi n}$ for wide range of DM masses with the advent of modern telescopes. The thermal emissions from old NSs in the infra-red and visible frequencies can be detectable in telescopes, such as the JWST. In particular, JWST aims to detect NSs with the red-shifted surface temperatures ($T^{\infty}_{S}$) within 2000 K to 40000 K Raj:2024kjq ; Chatterjee:2022dhp , which can determine the cooling procedure of old NSs. Therefore, one can test various late-time heating scenarios related to astrophysics of NS, including the DM annihilation effects. Here, we use the sensitivity of the JWST to probe the DM-neutron cross-section, assuming the late-time heating coming only from DM annihilations. The typical surface temperature emerges in case of DM annihilation is around $10^{-6}$ MeV ($10^{4}K$) for $T_{NS}=10^{-5}$ MeV, as apparent from Eq.13. In the both panels of Fig.6, the region above the orange dot-dashed line can be probed with the detection of thermal emissions of 10 Gyr old NS having $T^{\infty}_{S}\geq 2000$ K. To remind, the surface temperature of the NS is related to $T^{f}_{NS}$, which is dependent on $m_{\chi}C_{c}$ (see Eq.14) in the capture-annihilation equilibrium. Consequently, $\sigma_{\chi n}$ is determined with mild dependence on the DM mass, $m_{\chi}$ through $\sigma_{g}$ apparent from Eq.2.
@@ 4 Outlook
Our present study is about bosonic DM particles forming either a BEC or a non-BEC state inside a NS, depending on DM observables, namely $m_{\chi}$, $\sigma_{\chi n}$ and $\expectationvalue{\sigma v}$. One of the important implications of our analysis is that direct detection experiments remain relevant for the electroweak scale DM (see Fig.3). Indirect detection of DM puts an upper limit on the $\expectationvalue{\sigma v}$, whereas our work provides a lower limit - keeping the mass around the electroweak scale unconstrained as well (see Fig.7).
As is apparent that the constraints on DM parameters are sensitive to how BH forms from different thermal states of DM particles inside the NS. So, the question is whether we can observationally distinguish two kinds of BHs, namely BH originated from the BEC state and BH originated from a non-BEC state. In fact, the distinction between a BEC state BH and a non-BEC state BH hinges on the observation of the collapse time of BHs, as defined in Eq.22. Gravitational wave signals associated with binary systems of NSs and BHs can be used to infer about the collapse time, which is quite different for two kinds of BHs in consideration. In a binary NS population, the typical time delay to complete the merger event can be as large as $\mathcal{O}$(Gyr)Zevin:2022dbo , whereas the collapse time is rather small as shown in Eq.23. Consequently, some of binary NSs in a population can implode into binary BHs before merging into one another. Future gravitational wave detectors can measure the relative fraction of binary NSs and binary BHs in such a population, in turn inferring about the collapse time. For computational details see Ref.Singh:2022wvw . Therefore, with sufficient statistics of GW events in the foreseeable future, BHs with different origins in our scenario can be distinguishable.
Now, in the absence of BH formation from the host NS, the distinction between a BEC and a non-BEC DM core can be possible via probing the equation of state (EOS) of the NS Ellis:2018bkr ; Karkevandi:2021ygv ; Buras-Stubbs:2024don ; Shakeri:2022dwg . In particular, the EOS of a NS is modified differently for a BEC and a non-BEC DM core, as DM particles in the BEC state essentially are confined within a smaller spatial region with negligible pressure than the latter. This is reflected in two NS observables, namely the mass-radius relation and the tidal deformability, which are measurable in the gravitational wave spectrum LIGOScientific:2017vwq ; Annala:2017llu of binary NS systems as well as in radio observations of pulsars. In a NS binary system, tidal forces would be different for a BEC and a non-BEC DM core, which would be reflected in the tidal deformability parameter. However, the exact merger dynamics is rather involved, therefore left for elsewhere ghosh .